OpenStudy (anonymous):

what is x+3+y and 21x+8y=-24

OpenStudy (anonymous):

13y/21+13/7

OpenStudy (anonymous):

do you mean: x + 3 = y 21x + 8y = -24

OpenStudy (anonymous):

what does x and y equal

OpenStudy (anonymous):

Maybe this will help olve by Substitution : // Solve equation [1] for the variable x [1] x = y + 3 // Plug this in for variable x in equation [2] [2] 21•(y +3) + 8y = -24 [2] 29y = -87 // Solve equation [2] for the variable y [2] 29y = - 87 [2] y = - 3 // By now we know this much : x = y+3 y = -3 // Use the y value to solve for x x = (-3)+3 = 0 Solution : {x,y} = {0,-3}

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome